OpenStudy (anonymous):

Gino designed a garden for his rectangular backyard. The design he drew on paper is 10 inches long. The length of his backyard is 40 feet. If his drawing is to scale, and his backyard is 28 feet wide, how many inches wide is his drawing? Write a proportion using words. Then, use the proportion to calculate the drawing width. Show your work.

OpenStudy (anonymous):

the ratio of the length of the garden to the width of the garden is equal to the length of the drawing to the width of the drawing. 40/28 = 10/w 40w = 280 w(idth) = 7 inches

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Sure Thing !